OpenStudy (anonymous):

Find the simple interest percent to yield $100 interest in 5 years when $500 is invested.

OpenStudy (anonymous):

Simplest is that $100 is 20% of $500. If compounded yearly, then 100 = 500(1 + r/5)^5 - 500 r is the annual interest rate you are solving for

OpenStudy (anonymous):

\[S.I=\frac{ p*r*t }{100 }\] S.I=$100 P=500 $ t=5 years r=?

OpenStudy (anonymous):

Will you always divide every problem by 100?

OpenStudy (anonymous):

rate is percent.

OpenStudy (anonymous):

ohhh ok, got you

OpenStudy (anonymous):

Is it clear that "simple interest" meant "annual percentage rate"? Probably.

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

it should be \[\Large 100 = \frac{500*r*5}{100}\]

jimthompson5910 (jim_thompson5910):

notice how I typed in 5 and NOT 50

jimthompson5910 (jim_thompson5910):

\[\Large 100 = \frac{500*r*5}{100}\] \[\Large 100 = \frac{2500r}{100}\] \[\Large 100*100 = 2500r\] \[\Large 10000 = 2500r\] \[\Large 2500r = 10000\] \[\Large r = ???\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

correct